# Bayesian Inference General Procedures for A Single-subject Test Study

### v2

@@ 2 Bayesian Inference Procedures for Single-subject’s Abnormality Detection
In this section, we introduce a Bayesian inference procedures to compare a single-subject with a control group, under the assumption that the control group observations follow a skewed Student’s $t$ distribution. First, we should verify whether the assumption holds. For practical applications, we recommend employing a modified version of the general goodness-of-fit test (Chen and Balakrishnan, 1995) tailored explicitly for the skewed Student’s $t$ distribution in the Appendix A. From now on, we assume that the control group data follows a skewed Student’s $t$ distribution in the subsequent sections.
@@ 2.1 Bayesian Inference General Procedures for a Single-subject Test (BIGPAST)
Let $x^{*}$ denote the observation of a single-subject. Our goal is to test whether $x^{*}$ is a random sample from the same distribution as the control group. Let $f_{c}(\alpha,\nu,\xi,\omega)$ represent the skewed Student’s $t$ density function of $x_{1},x_{2},\ldots,x_{n}$. The hypothesis testing can be formulated as follows:
@@
where $\not\sim$ represents that $x^{*}$ is not a random sample from $f_{c}(\alpha,\nu,\xi,\omega)$. We propose the Bayesian inference general procedures for a single-subject test under the null hypothesis $\mathcal{H}_{0}$. The procedures are as follows:
@@
    Choose the proposed Jeffery prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ in Section 2.2 as the joint prior of $\alpha,\nu,\xi,\omega$.
@@
    Compute the posterior distribution of $\alpha,\nu,\xi,\omega$ based on the control group observations $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and the prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$. By Bayes’s Theorem, we have
@@
    Draw $m$ samples of $\alpha,\nu,\xi,\omega$ from the posterior distribution $\boldsymbol{\pi}(\alpha,\nu,\xi,\omega|\mathbf{x})$ by Metropolis-Hastings algorithm (Hastings, 1970) with normal proposal, denoted as $\alpha_{j},\nu_{j},\xi_{j},\omega_{j}$, $j=1,2,\ldots,m$.
@@
    For each $\alpha_{j},\nu_{j},\xi_{j},\omega_{j}$, draw $s$ random samples from the skewed Student’s $t$ distribution, denoted as $x_{1j},x_{2j},\ldots,x_{sj}$, $j=1,2,\ldots,m$. Sort the observations $x_{1j},x_{2j},\ldots,x_{sj}$, $j=1,2,\ldots,m$, in ascending order and label them as $x_{(1)},x_{(2)},\ldots,x_{(B)}$ where $B=ms$.
@@
Remarks
@@
When $\alpha\to+\infty$ or $\infty$- ∞, the skewed Student’s $t$ distribution degenerates to half Student’s $t$ distribution. Then, the two-sided credible interval in Step 5 above should be $[x_{(1)},x_{(\lceil B(1-\alpha^{*})\rceil)}]$ if $\alpha=+\infty$ or $[x_{(\lfloor B\alpha^{*}\rfloor)},x_{(B)}]$ if $\alpha=-\infty$. Furthermore, procedures 1-3 above can be extended to other distributions, provided that a prior can be assigned for a specific density distribution.
2.2 Jefferys Priors
@@
In Bayesian probability, the Jeffreys prior (Jeffreys, 1946) refers to a prior density function that is proportional to the square root of the determinant of the Fisher information matrix. In this section, we proposed an explicit Jeffery prior for the parameters of skewed Student’s $t$ distribution and compare it to the existing priors.
Let $f(z|\alpha,\nu)$ be the non-linear skewed Student’s $t$ density function (Azzalini and Capitanio, 2003), defined as follows:
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$ and $w=\alpha zr(z,\nu)$, $t(\cdot|\nu)$ represents the density function of $t$ distribution with degree of freedom $\nu$, and $T(\cdot|\nu+1)$ represents the cumulative density function of $t$ distribution with degree of freedom $\nu+1$. Let $\xi$ and $\omega$ be the location and scale parameters, with the substitution $z=(x-\xi)/\omega$ in Equation 1, then the skewed Student’s $t$ density function is given by
@@
In the literature, two types of priors are commonly used for $\xi$ and $\omega$: the partial information prior (Sun and Berger, 1998) and the independence Jefferys prior (Rubio and Steel, 2014). Both of these priors assume that the joint prior of $\xi$ and $\omega$, denoted as $\boldsymbol{\pi}(\xi,\omega)$, is proportional to $\omega^{-1}$. Consequently, the joint prior of $\alpha,\nu,\xi,\omega$ can be expressed as follows:
@@
The joint prior of $\alpha$ and $\nu$, $\boldsymbol{\pi}(\alpha,\nu)$, may be specified by the Jefferys prior (Branco et al., 2013) or the natural informative prior (Dette et al., 2018). Branco et al. (2013) proposed the following prior for general non-linear skewed Student’s $t$ density function:
@@
where $\boldsymbol{\pi}(\nu)$ is proposed by Fonseca et al. (2008) for the Student’s $t$ distribution and given by
$\boldsymbol{\pi}(\nu)\propto\left(\frac{\nu}{\nu+3}\right)^{1/2}\left\{\psi^{(1)}\left(\frac{\nu}{2}\right)-\psi^{(1)}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+3)}{\nu(\nu+1)^{2}}\right\}^{1/2},$
@@
Dette et al. (2018) proposed the so-called natural informative prior of the parameter $\alpha$ based on the total variance distance, which is given by
$BTV(\alpha|\theta_{1},\theta_{2})=\frac{1}{B(\theta_{1},\theta_{2})}\left(\frac{\tan^{-1}(\alpha)}{\pi}+\frac{1}{2}\right)^{\theta_{1}-1}\left(\frac{-\tan^{-1}(\alpha)}{\pi}+\frac{1}{2}\right)^{\theta_{2}-1}\frac{1}{\pi(1+\alpha^{2})},$
@@
A notable concern with $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$ is that the marginal prior of $\nu$ is independent of $\alpha$. However, given that the skewed Student’s $t$ distribution is a generalisation of the Student’s $t$ distribution, the marginal prior of $\nu$ should ideally depend on $\alpha$. To address this issue, we propose the Jefferys prior for the skewed Student’s $t$ density function, as detailed in Theorem 2.1.
@@
Let $f(z|\alpha,\nu)$ be the non-linear skewed Student’s $t$ density function, defined as follows:
@@
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$ and $w=\alpha zr(z,\nu)$. Given any $\nu>0$, let $\boldsymbol{\pi}^{J}(\alpha,\nu)$ represent the joint Jefferys prior for $\alpha$ and $\nu$, we have
@@
$I_{\nu\nu}$ $=\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]+\frac{1}{4}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]+\frac{1}{4}\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]+d^{2}_{\nu}$
$\qquad+\frac{1}{2}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]+d_{\nu}\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]$
$I_{\alpha\nu}$ $=-\frac{\pi\alpha\Gamma\left(\frac{\nu}{2}+1\right)^{2}\left((\nu+4)H_{5}-\frac{\left(\alpha^{2}(2\nu+3)+(\nu+3)\sigma_{\nu+1}^{2}\right)H_{6}}{\sigma_{\nu+1}^{2}}\right)}{8\left(\alpha^{2}+\sigma_{\nu+1}^{2}\right)\Gamma\left(\frac{\nu+1}{2}\right)\Gamma\left(\frac{\nu+5}{2}\right)},$
@@
$\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]$ $=\frac{\pi^{5/2}\Gamma\left(\frac{\nu}{2}+1\right)\left((\nu+3)H_{1}-(\nu+3)H_{2}-H_{3}+H_{4}\right)}{8\nu^{2}\Gamma\left(\frac{\nu+5}{2}\right)B\left(\frac{\nu}{2},\frac{1}{2}\right)B\left(\frac{\nu+1}{2},\frac{1}{2}\right)^{2}}.$
Furthermore, we define $d_{\nu}\coloneqq-\frac{1}{2}\psi\left(\frac{\nu}{2}\right)+\frac{1}{2}\psi\left(\frac{\nu+2}{2}\right)+2c_{\nu}*g_{1}(\nu)$, where $c_{\nu}$ is a constant that depends solely on $\nu$. The hypergeometric functions, denoted as $H_{1},H_{2},H_{3},H_{4},H_{5},H_{6}$, along with $h^{2}(w)$ and $g_{1}(\nu)$ are defined in the proof. The term $\sigma^{2}_{\nu+1}$ is defined in Lemma A.1. The functions $\psi(\cdot)$ and $\psi^{(1)}(\cdot)$ represent the digamma and trigamma function respectively.
The proof of Theorem 2.1 can be found in Appendix. Upon applying a Taylor expansion, we find that as $\alpha\rightarrow 0$, the order of $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ with respect to $\alpha$ is $O(1+\alpha^{2})$. Conversely, as $\alpha\rightarrow+\infty$, the order becomes $O(\alpha^{-3/2})$, aligning with the findings of Branco et al. (2013). The computation of $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ relies solely on the basic functions provided by the Python packages scipy and numpy. As such, the formulas in Theorem 2.1 can be directly implemented to define a function for computing Jefferys prior. Additionally, we have developed a Python package, skewt-scipy (https://pypi.org/project/skewt-scipy/), specifically for the skewed Student’s $t$ distribution.
For the Bayesian estimation of $\nu$, we anticipate that $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ will outperform $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$. Because $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ depends on $\alpha$, unlike $\boldsymbol{\pi}(\nu)$ in $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ is independent of $\alpha$. This is verified by the comparison study in Section 3. The Jefferys prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ is a suitable choice for the skewed Student’s $t$ density function. The BIGPAST based on Jefferys prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ is robust to detect the departures from normality and outperforms the existing approaches based on the assumption of normal distribution.

### v4

@@ 2 Bayesian Inference Procedures for Single-subject’s Abnormality Detection
In this section, we introduce Bayesian inference procedures to compare a single-subject with a control group, assuming that the control group observations follow a skewed Student $t$ distribution. First, we should verify whether the assumption holds. For practical applications, we recommend employing a modified version of the general goodness-of-fit test (Chen and Balakrishnan, 1995) tailored explicitly for the skewed Student $t$ distribution in the Appendix A. From now on, we assume that the control group data follows a skewed Student $t$ distribution in the subsequent sections.
@@ 2.1 Bayesian Inference General Procedures for a Single-subject Test (BIGPAST)
Let $x^{*}$ denote the observation of a single-subject. We aim to test whether $x^{*}$ is a random sample from the same distribution as the control group. Let $f_{c}(\alpha,\nu,\xi,\omega)$ represent the skewed Student $t$ density function of $x_{1},x_{2},\ldots,x_{n}$. The hypothesis testing can be formulated as follows:
@@
where $\not\sim$ represents that $x^{*}$ is not an observation from $f_{c}(\alpha,\nu,\xi,\omega)$. We propose the Bayesian inference general procedures for a single-subject test under the null hypothesis $\mathcal{H}_{0}$. The procedures are as follows:
@@
    Choose the proposed Jeffery prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ (defined below) as the joint prior of $\alpha,\nu,\xi,\omega$.
@@
    Compute the posterior distribution of $\alpha,\nu,\xi,\omega$ based on the control group sample $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and the prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$. By Bayes’s Theorem, we have
@@
    Draw $m$ samples of $\alpha,\nu,\xi,\omega$ from the posterior distribution $\boldsymbol{\pi}(\alpha,\nu,\xi,\omega|\mathbf{x})$ using the Metropolis-Hastings algorithm (Hastings, 1970)Algorithm 1 lists the details of sampling, the samples are denoted as $\alpha_{j},\nu_{j},\xi_{j},\omega_{j}$, $j=1,2,\ldots,m$.
@@
    For each $\alpha_{j},\nu_{j},\xi_{j},\omega_{j}$, draw $s$ random samples from the skewed Student $t$ distribution, denoted as $x_{1j},x_{2j},\ldots,x_{sj}$, $j=1,2,\ldots,m$. Sort the observations $x_{1j},x_{2j},\ldots,x_{sj}$, $j=1,2,\ldots,m$, in ascending order and label them as $x_{(1)},x_{(2)},\ldots,x_{(B)}$ where $B=ms$.
@@
Data: Initial parameters: $(\alpha_{0},\nu_{0},\xi_{0},\omega_{0})$, data: $\mathbf{x}$, size $m_{0}=10000$, burn-in: $b=0.4$, step size: $\delta=0.5$.
Result: Samples from the posterior distribution
1 $n_{0}\leftarrow 4$;
2 Θzeros matrix with $\Theta\leftarrow\text{zeros matrix with dimension:}(m_{0},n_{0})$;
3 $(\alpha^{\text{old}},\nu^{\text{old}},\xi^{\text{old}},\omega^{\text{old}})\leftarrow(\alpha_{0},\nu_{0},\xi_{0},\omega_{0})$;
4 $i\leftarrow 0$;
5 while $i<m_{0}$ do
6       $\nu^{\text{new}}\leftarrow\nu^{\text{old}}+\delta\Phi^{-1}[\Phi(-\nu^{\text{old}}/\delta)+u_{\nu}\Phi(\nu^{\text{old}}/\delta)]$;
7       $p_{\nu}\leftarrow\log\Phi(\nu^{\text{old}}/\delta)-\log\Phi(\nu^{\text{new}}/\delta)$;
8       $\omega^{\text{new}}\leftarrow\omega^{\text{old}}+\delta\Phi^{-1}[\Phi(-\omega^{\text{old}}/\delta)+u_{\omega}\Phi(\omega^{\text{old}}/\delta)]$;
9       $p_{\omega}\leftarrow\log\Phi(\omega^{\text{old}}/\delta)-\log\Phi(\omega^{\text{new}}/\delta)$;
10       $\alpha^{\text{new}}\leftarrow\alpha^{\text{old}}+\epsilon_{\alpha}$;
11       $\xi^{\text{new}}\leftarrow\xi^{\text{old}}+\epsilon_{\xi}$;
12       $p_{\text{new}}\leftarrow\log(\boldsymbol{\pi}(\alpha^{\text{new}},\nu^{\text{new}},\xi^{\text{new}},\omega^{\text{new}}|\mathbf{x}))$;
13       $p_{\text{old}}\leftarrow\log(\boldsymbol{\pi}(\alpha^{\text{old}},\nu^{\text{old}},\xi^{\text{old}},\omega^{\text{old}}|\mathbf{x}))$;
14       $A\leftarrow p_{\text{new}}-p_{\text{old}}+p_{\nu}+p_{\omega}$;
15       if $\log(u)\leq A$ then
16             $(\alpha^{\text{old}},\nu^{\text{old}},\xi^{\text{old}},\omega^{\text{old}})\leftarrow(\alpha^{\text{new}},\nu^{\text{new}},\xi^{\text{new}},\omega^{\text{new}})$;
17
18      $\Theta[i,:]\leftarrow(\alpha^{\text{old}},\nu^{\text{old}},\xi^{\text{old}},\omega^{\text{old}})$;
19       $i\leftarrow i+1$;
20
21return $\Theta[\lfloor m_{0}*b\rfloor+1:,:]$;
Algorithm 1 Metropolis-Hastings Algorithm
In Algorithm 1, the Metropolis-Hastings algorithm draws samples from the posterior distribution. The step size $\delta$ is set to 0.5, and the burn-in rate $b$ is set to 0.4. The algorithm returns the samples after the burn-in period. $u_{\nu},u_{\omega},u$ are random numbers drawn from the uniform distribution $U(0,1)$. $\epsilon_{\alpha}$ and $\epsilon_{\xi}$ are drawn from the Normal distribution with mean zero and standard deviation $\delta$. As $\nu$ and $\omega$ are positive, the proposal distributions for $\nu$ and $\omega$ are the truncated normal distributions with mean $\nu^{\text{old}}$, $\omega^{\text{old}}$ and standard deviation $\delta$, respectively, see lines 6-9 of Algorithm 1. The proposal distributions for $\alpha$ and $\xi$ are normal distributions with mean $\alpha^{\text{old}}$, $\xi^{\text{old}}$ and standard deviation $\delta$, respectively, see lines 10-11 of Algorithm 1. The truncated proposal distributions do not have symmetric densities, so we have to adjust the acceptance rate $A$ by adding the logarithm of the ratios, i.e., $p_{\nu}$ and $p_{\omega}$ to the acceptance rate $A$, see lines 12-14 of Algorithm 1.
Steps 3, 4 and 5 employ the nested sampling method to determine the critical interval. One popular way is to use the following steps $3^{\prime}$ and $4^{\prime}$:
  1. 3
  2. 4
    Given the significance level $\beta$, the credible interval of $x^{*}$ given $\mathbf{x}$ is $[x^{\mathrm{MAP}}_{\beta/2},x^{\mathrm{MAP}}_{1-\beta/2}]$, where $x^{\mathrm{MAP}}_{\beta/2}$ and $x^{\mathrm{MAP}}_{1-\beta/2}$ are the $\beta/2$ and $21-\beta/$ 2 quantiles of $\boldsymbol{\pi}(\hat{\alpha}_{\mathrm{MAP}},\hat{\nu}_{\mathrm{MAP}},\hat{\xi}_{\mathrm{MAP}},\hat{\omega}_{\mathrm{MAP}}|\mathbf{x})$, respectively. If $x^{*}$ does not fall within this credible interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\beta$.
However, in practice, we found that the maximum posterior estimation is time-consuming, and the steps $3^{\prime}$ and $4^{\prime}$ do not perform better than the steps 3, 4 and 5, see the comparison study in Section 3.4. Therefore, we recommend using the nested sampling method to determine the critical interval.
@@ 19       $i\leftarrow i+1$;
Moreover, when $\alpha\to+\infty$ or $\infty$- ∞, the skewed Student $t$ distribution degenerates to half Student $t$ distribution. Then, the two-sided credible interval in Step 5 above should be slightly modified as $[x_{(1)},x_{(\lceil B(1-\beta)\rceil)}]$ if $\alpha\to+\infty$ or $[x_{(\lfloor B\beta\rfloor)},x_{(B)}]$ if $\alpha\to-\infty$.
2.2 Jeffreys Priors
@@
In Bayesian probability, Jeffreys prior (Jeffreys, 1946) refers to a prior density function proportional to the square root of the determinant of the Fisher information matrix. In this section, we proposed an explicit Jeffery prior for the parameters of the skewed Student $t$ distribution and compared it to existing priors.
@@
In the literature, two types of priors are commonly used for $\xi$ and $\omega$ in Equation 2: the partial information prior (Sun and Berger, 1998) and the independence Jeffreys prior (Rubio and Steel, 2014). Both of these priors assume that the joint prior of $\xi$ and $\omega$, denoted as $\boldsymbol{\pi}(\xi,\omega)$, is proportional to $\omega^{-1}$. Consequently, the joint prior of $\alpha,\nu,\xi,\omega$ can be expressed as follows:
@@
The joint prior of $\alpha$ and $\nu$, $\boldsymbol{\pi}(\alpha,\nu)$, may be specified by the Jeffreys prior (Branco et al., 2013) or the natural informative prior (Dette et al., 2018).  Branco et al. (2013) proposed the following prior for the general skewed Student $t$ density function:
@@
where $\boldsymbol{\pi}(\nu)$ is proposed by Fonseca et al. (2008) for the Student $t$ distribution and given by
${\footnotesize\boldsymbol{\pi}(\nu)\propto\left(\frac{\nu}{\nu+3}\right)^{1/2}\left\{\psi^{(1)}\left(\frac{\nu}{2}\right)-\psi^{(1)}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+3)}{\nu(\nu+1)^{2}}\right\}^{1/2}},$
@@
Dette et al. (2018) proposed the so-called natural informative prior of the parameter $\alpha$ based on the total variation distance, which is given by
$BTV(\alpha|\theta_{1},\theta_{2})=\frac{1}{B(\theta_{1},\theta_{2})}\left(\frac{\tan^{-1}(\alpha)}{\pi}+\frac{1}{2}\right)^{\theta_{1}-1}\left(\frac{-\tan^{-1}(\alpha)}{\pi}+\frac{1}{2}\right)^{\theta_{2}-1}\frac{1}{\pi(1+\alpha^{2})}$
@@
A notable concern with $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$ is that the marginal prior of $\nu$ is independent of $\alpha$. However, given that the skewed Student $t$ distribution is a generalisation of the Student $t$ distribution, the marginal prior of $\nu$ should ideally depend on $\alpha$. We propose the Jeffreys prior for the skewed Student $t$ density function, as detailed in Theorem 2.1.
@@
Let $f(z|\alpha,\nu)$ be the skewed Student $t$ density function, defined as follows:
@@
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$ and $w=\alpha zr(z,\nu)$. Given any $\nu>0$, let $\boldsymbol{\pi}^{J}(\alpha,\nu)$ represent the joint Jeffreys prior for $\alpha$ and $\nu$. We have
@@
$I_{\nu\nu}$ $=\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]+\frac{1}{4}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]+\frac{1}{4}\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]$
$\quad+d^{2}_{\nu}+d_{\nu}\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]+\frac{1}{2}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]$
$I_{\alpha\nu}$ $=-\frac{\pi\alpha\Gamma\left(\frac{\nu}{2}+1\right)^{2}}{8\left(\alpha^{2}+\sigma_{\nu+1}^{2}\right)\Gamma\left(\frac{\nu+1}{2}\right)\Gamma\left(\frac{\nu+5}{2}\right)}\left((\nu+4)H_{5}-\frac{\left(\alpha^{2}(2\nu+3)+(\nu+3)\sigma_{\nu+1}^{2}\right)H_{6}}{\sigma_{\nu+1}^{2}}\right),$
@@
$\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]=\psi\left(\frac{\nu}{2}\right)-\psi\left(\frac{\nu+1}{2}\right),\quad\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]=-\frac{1}{\nu^{2}+\nu}$
$\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]=\frac{1}{\nu^{2}+3\nu},\quad\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]=\left(\psi\left(\frac{\nu}{2}\right)-\psi\left(\frac{\nu+1}{2}\right)\right)^{2}\psi^{(1)}\left(\frac{\nu}{2}\right)-\psi^{(1)}\left(\frac{\nu+1}{2}\right)$
$\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]=\frac{\pi^{5/2}\Gamma\left(\frac{\nu}{2}+1\right)}{8\nu^{2}\Gamma\left(\frac{\nu+5}{2}\right)B\left(\frac{\nu}{2},\frac{1}{2}\right)B\left(\frac{\nu+1}{2},\frac{1}{2}\right)^{2}}\left((\nu+3)H_{1}-(\nu+3)H_{2}-H_{3}+H_{4}\right).$
Furthermore, we define $d_{\nu}\coloneqq-\frac{1}{2}\psi\left(\frac{\nu}{2}\right)+\frac{1}{2}\psi\left(\frac{\nu+2}{2}\right)+2c_{\nu}*g_{1}(\nu)$, where $c_{\nu}$ is a constant that depends solely on $\nu$. The hypergeometric functions, denoted as $H_{1},H_{2},H_{3},H_{4},H_{5},H_{6}$, along with $h^{2}(w)$ and $g_{1}(\nu)$ are defined in the proof. ${}_{2}F_{1}\left(a,b;c;d\right)$ represents the hypergeometric function. The term $\sigma^{2}_{\nu+1}$ is defined in Lemma A.1. The functions $\psi(\cdot)$ and $\psi^{(1)}(\cdot)$ represent the digamma and trigamma function respectively.
The proof of Theorem 2.1 can be found in Appendix. Upon applying a Taylor expansion, we find that as $\alpha\rightarrow 0$, the order of $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ with respect to $\alpha$ is $O(1+\alpha^{2})$. Conversely, as $\alpha\rightarrow+\infty$, the order becomes $O(\alpha^{-3/2})$, aligning with the findings of Branco et al. (2013). The computation of $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ relies solely on the basic functions provided by the Python packages scipy and numpy. As such, the formulas in Theorem 2.1 can be directly implemented to define a function for computing the Jeffreys prior. Additionally, we have developed a Python package, skewt-scipy (https://pypi.org/project/skewt-scipy/), for the skewed Student $t$ distribution.
For the Bayesian estimation of $\nu$, we anticipate that $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ will outperform $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$. Because $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ depends on $\alpha$, unlike $\boldsymbol{\pi}(\nu)$ in $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ is independent of $\alpha$. This is verified by the comparison in Section 3.1. The Jeffreys prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ is a suitable choice for the skewed Student $t$ density function, see Table 2Table 2 also shows that the BIGPAST based on the Jeffreys prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ is robust to detecting the departures from normality and outperforms the existing approaches based on the assumption that the data come from a Normal distribution.